OpenStudy (anonymous):

How would I go about solving for x in this equation? 49^x = 7^(x – 3)

OpenStudy (anonymous):

\[49^{x}=7^{x-3}\]

OpenStudy (johnweldon1993):

Hint* \[\large a^x = x \times \ln a\] So take the log of both sides here... \[\large \ln 49^x = \ln 7^{x - 3}\] becomes \[\large x \times \ln 49 = (x-3) \times \ln 7 \] Can you take it from there?

OpenStudy (anonymous):

yes thank you so much!!!:)

OpenStudy (johnweldon1993):

Anytime :)

jhonyy9 (jhonyy9):

@cayssaday do you understand it ? but this is possible solving without logarithm natural so in the left side there is 49 on exponent x ,yes ? but you know that 49 is equal 7 squared so in this case in the left side will be 7^(2x) and on the right side there is 7^(x-3) so than now you see that there is 7 on exponent 2x equal 7 on exponent (x-3) so you know that than the numbers ,like in this case 7 ,with one exponent are the same so than the left side will be equal right side than the exponents are equales so make equale 2x=x-3 solve it for x and will get the right value for x hope this is understandably sure right easy

OpenStudy (anonymous):

@jhonyy9 Thank yous so much, your explanation was awesome and helped a lot!!:)

jhonyy9 (jhonyy9):

was my pleasure good luck